OpenStudy (anonymous):

What is the exact diameter of an area rug with a circumference of forty-two feet?

OpenStudy (anonymous):

What is the exact diameter of an area rug with a circumference of forty-two feet? Answer (the diameter is equal to two radii) 2 pi r= 42 d pi = 42 diameter= 42/pi

OpenStudy (anonymous):

so what would ur answer be???

OpenStudy (anonymous):

ANS: forty-two divided by pi feet

OpenStudy (anonymous):

Circumference = pi * diameter 42 = 3.14 * d Solve for d. Divide both sides by 3.14 42 /3.14 = d

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

thanks guys!

OpenStudy (anonymous):

think u can help me im behinde in school work my parents forgot to pay the internet bill http://openstudy.com/study#/updates/513e1719e4b029b0182bdd80